OpenStudy (anonymous):

Aruna cut a cake into two halves and cuts one half into smaller pieces of equal size. Each of the small pieces is twenty grams in weight. If she has seven pieces of the cake in all with her, how heavy was the original cake.

OpenStudy (anonymous):

hey plz help

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

explain this to me with fully clear steps

OpenStudy (anonymous):

with full statements.

OpenStudy (paxpolaris):

 ok first we cut the cake in half

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

dude just show me full steps. no need diagrams

OpenStudy (anonymous):

detailed ans with statements.

OpenStudy (paxpolaris):

next we need to cut one of the halves into equal pieces so that we have a total of 7 pieces... So, how many pieces do we need to cut that half into??? (i need the diagram to think)

OpenStudy (anonymous):

six

OpenStudy (anonymous):

dude with statements

OpenStudy (anonymous):

dude i know that the ans is 240g but i need u to show me all the steps with statements

OpenStudy (paxpolaris):

 correct

OpenStudy (anonymous):

next ?

OpenStudy (paxpolaris):

so the small piece is 1/6th of 1/2 of the cake .... or 1/12th

OpenStudy (anonymous):

ok

OpenStudy (paxpolaris):

so we need to do 20 times twelve

OpenStudy (anonymous):

now show me the full detailed ans with steps

OpenStudy (anonymous):

without diagrams

OpenStudy (anonymous):

with statements too

OpenStudy (paxpolaris):

Aruna cut a cake into two halves. Then she cut one of the halves into 6 equal pieces to get a total of 7 pieces. Each small piece is \(\frac16 \times \frac12= \frac 1 {12}\) of the whole. If \(\frac 1 {12}\)th of the cake weighs 20 g, The whole cake will weigh \(20 \times 12= 240\) g